Two point charge Q1 and Q2 are 3m apart, and their combined charge is 20micro coulumb.

(a) if one repels the other with a force of 0.075N, what are the two charges?

(b) if one attract the other with a force of 0.525N, what are the magnitude of the charges?


1
Expert's answer
2022-01-16T13:10:27-0500

"q_1q_2=\frac{Fr^2}k,"

a)

"q_1q_2=75,"

"q_1+q_2=20,"

"q_1=15~\mu C," "q_2=5~\mu C,"

b)

"q_1q_2=25,"

"q_1+q_2=20,"

"q_1=18.7~\mu C," "q_2=1.3~\mu C."